OpenStudy (anonymous):

Find the local and absolute extreme values of the function on the given interval. f(x)=x/(2x^2+5), [0,2]

OpenStudy (anonymous):

you got your derivative? what's f '(x) =

OpenStudy (anonymous):

I need help with the quotient rule was this the one with the special thing again?

OpenStudy (anonymous):

special thing?

OpenStudy (anonymous):

hi dee hi hi dee lo thing?

OpenStudy (anonymous):

if \(\huge f(x)=\frac{u}{v} \) then \(\huge f'(x)=\frac{v\cdot u'-u \cdot v'}{v^2} \)

OpenStudy (anonymous):

oh yeah... "lo d hi minus hi d lo all over lolo"

OpenStudy (anonymous):

so let u=x, v=2x^2+5... so u' = 1, v' = 4x now plug it into that formula for the quotient rule...:)

OpenStudy (anonymous):

(2x^2)(1)-(x)(4x)/(2x^2+5)^2

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

\[\huge f'(x)=\frac{v\cdot u' -u \cdot v' }{v^2} \] \[\huge f'(x)=\frac{(2x^2+5)\cdot (1) -(x) \cdot (4x) }{(2x^2+5)^2} \]

OpenStudy (anonymous):

Yeah thats what i got!

OpenStudy (anonymous):

oh... i thought you wrote something else....

OpenStudy (anonymous):

the answer you wrote didn't have (2x^2+5) in the numerator.

OpenStudy (anonymous):

whoops, that is what i wrote down though

OpenStudy (anonymous):

where do i go from there?

OpenStudy (anonymous):

simplify that numerator... and set it equal to zero and solve...

OpenStudy (anonymous):

\[\huge f'(x)=0\rightarrow\frac{5-2x^2 }{(2x^2+5)^2}=0\] \[\huge 5-2x^2 =0\]

OpenStudy (anonymous):

Thats not what i got

OpenStudy (anonymous):

did i mess up on the simplification of the numerator?

OpenStudy (anonymous):

oh ok i just did it wrong

OpenStudy (anonymous):

so solving that will give you two values, but only 1 of them is in the interval [0, 2]

OpenStudy (anonymous):

how did you get two values, i only got one i got x=1.5811

OpenStudy (anonymous):

is it 1.6 and -1.6?

OpenStudy (anonymous):

\[\huge 5-2x^2 =0\] \[\huge 5=2x^2\] \[\huge \frac{5}{2}=x^2\]

OpenStudy (anonymous):

\[\huge \frac{5}{2}=x^2\] \[\huge \sqrt{\frac{5}{2}}=x^2\]

OpenStudy (anonymous):

sorry... i got the aw snap message...

OpenStudy (anonymous):

instead of putting your answer as 1.6, i think it's better to keep it in exact form so at least you don't lose any accuracy.

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

yeah sorry

OpenStudy (anonymous):

but what do I do? is that right? 1.581139?

OpenStudy (anonymous):

yes... is it ok to put your answers in approximate decimal form instead of the exact form?

OpenStudy (anonymous):

I guess. I haven't really asked

OpenStudy (anonymous):

but is that the final answer, we dont have to do anything else?

OpenStudy (anonymous):

no... remember the question asks for local and global extrema...

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

all we did is find where f'(x) = 0....

OpenStudy (anonymous):

oh ok :) just kidding

OpenStudy (anonymous):

if you don't mind, i'll use the exact form of our answer... \(\large \sqrt{\frac{5}{2}}=\frac{\sqrt{10}}{2} \)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok... so we need to calculate all of these \[\huge f(0)= \] \[\huge f(2)= \] \[\huge f(\frac{\sqrt{10}}{2})= \]

OpenStudy (anonymous):

the first two is just the function evaluated at the endpoints of the interval you're looking at...

OpenStudy (anonymous):

uh huh

OpenStudy (anonymous):

f(0)= 0

OpenStudy (anonymous):

after you have those y-values, you can determine which are local/global max/mins.

OpenStudy (anonymous):

f(2)= 2/13

OpenStudy (anonymous):

correct..., correct.

OpenStudy (anonymous):

umm and im not sure how to figure out the last one

OpenStudy (anonymous):

\[(\sqrt{10}/2 )^{2}\] what does that equal

OpenStudy (anonymous):

\[\huge f(x)=\frac{x}{2x^2+5}\rightarrow \frac{\frac{\sqrt{10}}{2}}{2(\frac{\sqrt{10}}{2})^2+5} \] \[\huge \frac{\frac{\sqrt{10}}{2}}{2(\frac{\sqrt{10}}{2})^2+5} =\frac{\frac{\sqrt{10}}{2}}{2(\frac{10}{4})+5}=\frac{\frac{\sqrt{10}}{2}}{5+5} \]

OpenStudy (anonymous):

oh ok!

OpenStudy (anonymous):

\[\huge \frac{\frac{\sqrt{10}}{2}}{5+5}\rightarrow \frac{\frac{\sqrt{10}}{2}}{10}\rightarrow \frac{\sqrt{10}}{20} \]

OpenStudy (anonymous):

so what do you have as far as local/global max/min in your interval [0,2] ?

OpenStudy (anonymous):

umm im not sure.. whats the difference in the two

OpenStudy (anonymous):

a global extrema is the highest or lowest within the interval... a local extrema is only the highest or lowest within a "neighborhood"...

OpenStudy (anonymous):

so in other words, all global extremas ARE local extremas... but not the other way around...

OpenStudy (anonymous):

oh so would the local minimum would be 0?

OpenStudy (anonymous):

yes.. x=0 is a local minimum... BUT notice it is the lowest y-value in that interval so it is your global minimum...

OpenStudy (anonymous):

keep going... analyze the other two...

OpenStudy (anonymous):

what's the highest y-value?

OpenStudy (anonymous):

so the global max is \[\sqrt{10}\div20\]

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

what about at x=2?

OpenStudy (anonymous):

which would that be?

OpenStudy (anonymous):

i'm asking you that...

OpenStudy (anonymous):

im not sure that

OpenStudy (anonymous):

because we still have to find the local min and max and there is only on number left

OpenStudy (anonymous):

here's an easy "cheat" way of doing it... look at the graph of the points we have... if we were to connect them with a smooth curve what can you tell me?